OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (1, -3) and (3, -5). -1 1 0 one half

OpenStudy (anonymous):

@myininaya

OpenStudy (danjs):

(x,y) = ( 1 ,-3) and (x1,y1) = (3,-5) find \[slope = \frac{ y - y1 }{ x-x1 }\]

OpenStudy (anonymous):

right two , down two

OpenStudy (anonymous):



OpenStudy (anonymous):

ok @satellite73

OpenStudy (anonymous):

@DanJS not understanding

OpenStudy (danjs):

You find the difference in the 2 y values, and divide that by difference in the 2 x values

OpenStudy (anonymous):

right two, down two \[\huge \frac{-2}{2}\]

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

right @DanJS

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

yes it is \(-1\)

OpenStudy (anonymous):

ok so what is after that

OpenStudy (danjs):

\[\frac{ -5 +3 }{ 3 - 1 } = \frac{ -2 }{ 2 } = -1\]

OpenStudy (danjs):

you match -1 to the letter in the answers

OpenStudy (anonymous):

that is the answer

OpenStudy (anonymous):

@satellite73 @DanJS

OpenStudy (danjs):

?

OpenStudy (anonymous):

is that the answer @DanJS -1

OpenStudy (danjs):

Choose the correct slope of the line that passes through the points (1, -3) and (3, -5). "The slope is -1"

OpenStudy (anonymous):

thank you may you help me with another

OpenStudy (danjs):

new thread

OpenStudy (anonymous):

ok